I have a very interesting question for you my dear friends.


Question - Two persons decided to meet at a place and they decided their meeting time between 10 a.m. to 11 a.m.


But they put a condition that whosever will arrive first he will wait only for 10 minute for another person otherwise he will go.


What is the probabaility for their meeting?


Reply with appropriate solution.


Thanks

Grade:11

2 Answers

Chetan Mandayam Nayakar
11 years ago

answer=(1/6)+(1/6)+(2/6)=2/3

Anshul Mittal
11 years ago

can you please explain how??????